For the following set of cash flows,

   

Year Cash Flow
0 –$8,200           
1 4,800           
2 5,800           
3 3,900           

  

Required:
(a) What is the NPV at a discount rate of 0 percent?
(Click to select)$6,489$6,300.00$5,985$6,615$6,111

   

(b) What is the NPV at a discount rate of 12 percent?
(Click to select)$3,380.82$3,659.65$3,311.11$3,485.38$3,589.94

   

(c) What is the NPV at a discount rate of 22 percent?
(Click to select)$1,714.55$1,832.35$1,690.04$1,867.93$1,778.99
(d) What is the NPV at a discount rate of 27 percent?
(Click to select)$1,047.09$1,079.48$1,025.5$1,111.86$1,133.45
Answer #1

NPV = Present value of cash inflows - present value of cash outflows

a)

NPV = -8200 + 4,800 / ( 1 + 0)1 + 5,800 / ( 1 + 0)2 + 3,900 / ( 1 + 0)3

NPV = $6,300.00

b)

NPV = -8200 + 4,800 / ( 1 + 0.12)1 + 5,800 / ( 1 + 0.12)2 + 3,900 / ( 1 + 0.12)3

NPV = $3,485.38

c)

NPV = -8200 + 4,800 / ( 1 + 0.22)1 + 5,800 / ( 1 + 0.22)2 + 3,900 / ( 1 + 0.22)3

NPV = $1,778.99

d)

NPV = -8200 + 4,800 / ( 1 + 0.27)1 + 5,800 / ( 1 + 0.27)2 + 3,900 / ( 1 + 0.27)3

NPV = $1,079.48